how do i calculate the assiciated CaO from a dolomite value
1
Expert's answer
2012-11-16T02:24:38-0500
CaMg(CO3)2- dolomite
w(CaO) = M(CaO)/M(CaMg(CO3)2) x 100 %
M(CaMg(CO3)2) = 184 g/mol
M(CaO) = 56 g/mol
w(CaO) = 56/184 x 100% = 30.43 %